# Towards a multi-dimensional picture of nucleation: The example of liquid condensation through theory and rare event sampling

### v2

@@ Abstract
With two novel ideas, our work revisits the well-studied problem of liquid condensation from vapor. On the one hand, we use an innovative numerical strategy based on rare event sampling techniques to precisely measure structural features in atomistic simulations. On the other hand, an accessible theoretical framework assuming a multi-dimensional picture of nucleation is built to predict structural features of the critical nucleation cluster along with the associated free energy barrier and nucleation rate. Altogether, we show that the multi-dimensional picture of nucleation outperforms both qualitatively and quantitatively CNT and more recent theories, paving the way for a better understanding of nucleation.
@@
Theory - Our theoretical framework builds on the multi-variable nucleation theory [55, 56]. To simplify its application, we extend the capillary approximation in CNT by explicitly incorporating both the size $R$ and the density $\rho$ of the nucleating cluster. Consequently, the work of formation is given by $\Delta\varOmega(R,\rho)=-4\mathrm{\pi}{R}^{3}g_{n}/3+4\mathrm{\pi}{R}^{2}\gamma$. Here, $g_{n}$ is the driving force for nucleation, defined as $g_{n}=\mathcal{F}(\rho_{0})+\frac{\partial\mathcal{F}}{\partial\rho}\Big{|}_{\rho=\rho_{0}}(\rho-\rho_{0})-\mathcal{F}(\rho)$, with $\mathcal{F}$ the Helmholtz free energy per unit volume [23]. The surface tension of vapor-liquid interface, $\gamma$, is inspired by the Square Gradient Approximation (SGA), $\gamma=\kappa(\rho-\rho_{0})^{2}$, where $\kappa$ is assumed density-independent and can be computed from the surface tension at equilibrium [47, 57, 58, 25, 59]. For simplification purpose, we chose to ignore the dependence of surface tension with size [21]. We note that unlike in CNT, both $g_{n}$, and $\gamma$ are now treated as functions of the cluster density, $\rho$.
@@
Results - To begin, from Fig. 1(a) which shows representative critical clusters obtained with our rare-event sampling strategy, one can already observe that along with the size, the density of the nucleating cluster seem to decrease when increasing supersaturation. Then, more quantitatively, Fig. 1 (b-c) show the critical cluster radii and densities averaged over the whole distribution of clusters obtained in simulations. As predicted by other multi-dimensional nucleation theories [76, 47, 57], our numerical results confirm that the density of the critical cluster decreases with supersaturation. From the theoretical viewpoint, on the one hand, the critical cluster size is similarly predicted by all models. At the lowest supersaturation, there is a strong agreement with the simulation data. However, all models slightly underestimate the cluster size at intermediate supersaturation levels. On the other hand, regarding the critical cluster density, CNT fails at capturing the observed densification. Meanwhile, although it is able to qualitatively retrieve this decrease, the SGA diffuse-interface model [23, 60, 27, 29] overestimates the cluster density and exhibits a maximum that is not seen in simulations. In contrast, our two-variable model successfully reproduces the numerical results both qualitatively and quantitatively.
@@
Next, we use our theoretical model and CNT to evaluate the steady-state nucleation rate [See Fig. 2]. Computing the nucleation rate within the SGA model and comparing it with our two-variable theory would be very valuable since both approaches go beyond the simplistic one-variable CNT. However, this calculation is very difficult [29, 28] and not performed here. Then, we compare those predictions to numerical results obtained in the literature or using our own brute-force simulations reported in Table 1 in Appendix and analyzed with the Yasuoka-Matsumoto (YM) threshold method [33, 19]. Remarkably, the two-variable model performs very well at high supersaturation, where CNT is known to fail by several orders of magnitude. Our model contains no parameters adjusted to nucleation simulations, which demonstrates its predictive capability when properly applied and coupled to quantitative thermodynamics and kinetics data. Furthermore, contrary to CNT, the nucleation rate shows a maximum near the spinodal, an effect also observed in diffuse-interface models [29].
Figure 2: Evolution of the steady-state nucleation rate against initial density at $T=0.8\epsilon/k_{\mathrm{B}}$. Our brute-force simulation results are represented by blue up triangles. Results from Diemand et al. [19] are represented by green down triangles. Predictions from our two-variable nucleation theory and CNT are represented by a red solid line and a yellow dotted lines, respectively.
@@
Prompted by the confirmations of its quality, our model can be used to predict the work of formation, see Fig. 3. At low supersaturation, the three models agree because critical cluster density matches closely the equilibrium liquid density. In the intermediate supersaturation region, both our model and SGA predict much smaller nucleation barriers. It is counter-intuitive at first since the driving force for nucleating a dilute cluster is smaller as compared to a cluster with the equilibrium density, but it is balanced by a much lower surface energy cost. Finally, when approaching the spinodal limit, the two-variable nucleation theory recovers results consistent with diffuse-interface models: the density of critical cluster approaches the initial phase density, its size diverges, and the associated work of formation vanishes. This behavior provides a smooth transition between nucleation and spinodal decomposition, where the energy barrier is zero, as established in diffuse-interface models [23] and recovered in previous nucleation models with two parameters [77, 78, 79, 48, 80, 57, 58, 81, 25, 76, 47]. CNT, however, fails to capture this behavior, highlighting the improved physical accuracy of the two-variable model.
Figure 4: Contour plot of the work of formation overlaid with growth trajectories of critical clusters from simulations at $T=0.8\epsilon/k_{\mathrm{B}}$ and $\rho_{0}=0.03\sigma^{-3}$. The work of formation is computed using our two-variable nucleation theory, with the saddle point marked by a red hexagram. The red dash line and blue dotted line indicate the predicted nucleation flux direction and steepest decent direction at the saddle point, respectively. Growth trajectories from four independent simulations are represented by colored crosses, with corresponding mean values shown as dots. Error bars denote standard deviations. $(R,\rho)$ coordinates from theory and simulation are normalized by their corresponding critical values.
@@ Appendix E Incorporation of dynamical density functional theory (DDFT) into nucleation theory
Substituting this expression into Eq. (7), pre-multiplying by $\frac{\partial n}{\partial\boldsymbol{x}}$, and integrating over $r$ from 00 to infinity, we obtain
@@
$\mathbf{D}=D_{\mathrm{diff}}\begin{pmatrix}4\mathrm{\pi}R^{3}\dfrac{(\rho-\rho_{0})^{2}}{\rho_{0}}&4\mathrm{\pi}R^{4}\dfrac{\rho-\rho_{0}}{3\rho_{0}}\\ 4\mathrm{\pi}R^{4}\dfrac{\rho-\rho_{0}}{3\rho_{0}}&\frac{4}{9}\mathrm{\pi}R^{5}\big{(}\frac{1}{5\rho}+\frac{1}{\rho_{0}}\big{)}\end{pmatrix}.$ (10)

### v3

@@ Abstract
With two novel ideas, our work revisits the well-studied problem of liquid condensation from vapor. On the one hand, we use an innovative numerical strategy based on rare event sampling techniques to precisely measure structural features of critical clusters in atomistic simulations over a large range of supersaturation. On the other hand, an accessible theoretical framework assuming a multi-dimensional picture of nucleation is built to predict structural features of the critical nucleation cluster along with the associated free energy barrier and nucleation rate. Altogether, we show that the multi-dimensional picture of nucleation outperforms both qualitatively and quantitatively CNT and more recent theories, paving the way for a better understanding of nucleation.
@@
Theory - Our theoretical framework builds on the multi-variable nucleation theory [55, 56]. To simplify its application, we extend the capillary approximation in CNT by explicitly incorporating both the size $R$ and the density $\rho$ of the nucleating cluster. Consequently, the work of formation is given by $\Delta\varOmega(R,\rho)=-4\mathrm{\pi}{R}^{3}g_{n}/3+4\mathrm{\pi}{R}^{2}\gamma$. Here, $g_{n}$ is the driving force for nucleation, defined as $g_{n}=\mathcal{F}(\rho_{0})+\frac{\partial\mathcal{F}}{\partial\rho}\Big{|}_{\rho=\rho_{0}}(\rho-\rho_{0})-\mathcal{F}(\rho)$, with $\mathcal{F}$ the Helmholtz free energy per unit volume [23]. The surface tension of vapor-liquid interface, $\gamma$, is inspired by the Square Gradient Approximation (SGA), $\gamma=\kappa(\rho-\rho_{0})^{2}$, where $\kappa$ is assumed density-independent and can be computed from the surface tension at equilibrium [47, 57, 58, 25, 59]. At first we ignore the dependence of surface tension with size [21]. We note that unlike in CNT, both $g_{n}$, and $\gamma$ are now treated as functions of the cluster density, $\rho$.
@@
Results - To begin, from Fig. 1(a) which shows representative critical clusters obtained with our rare-event sampling strategy, one can already observe that along with the size, the density of the nucleating cluster seem to decrease when increasing supersaturation. Then, more quantitatively, Fig. 1 (b-c) show the critical cluster radii and densities averaged over the whole distribution of clusters obtained in simulations. As predicted by other multi-dimensional nucleation theories [76, 47, 57], our numerical results confirm that the density of the critical cluster decreases with supersaturation. From the theoretical viewpoint, on the one hand, the critical cluster size is similarly predicted by all models. At the lowest supersaturation, there is a strong agreement with the simulation data. However, all models slightly underestimate the cluster size at intermediate supersaturation levels. On the other hand, regarding the critical cluster density, CNT fails at capturing the observed densification. Meanwhile, although it is able to qualitatively retrieve this decrease, the SGA diffuse-interface model [23, 75, 27, 29] overestimates the cluster density and exhibits a maximum that is not seen in simulations. In contrast, our two-variable model successfully reproduces numerical results both qualitatively and quantitatively.
@@
Next, we use our theoretical model and CNT to evaluate the steady-state nucleation rate [See Fig. 2]. Computing the nucleation rate within the SGA model and comparing it with our two-variable theory would be very valuable since both approaches go beyond the simplistic one-variable CNT. However, coupling the SGA model with the DDFT dynamics to calculate the nucleation rate is not trivial [29, 28] due to the high dimensionality of diffuse-interface formalisms. This calculation is not performed here. Reduction of the problem to two dimensions in our sharp-interface theory, but with the cluster density allowed to vary, results in a straightforward calculation of the nucleation rate. Then, we compare those predictions to numerical results obtained in the literature or using our own brute-force simulations reported in Table 1 in Appendix and analyzed with the Yasuoka-Matsumoto (YM) threshold method [33, 19]. Remarkably, the two-variable model performs very well at high supersaturation, where CNT is known to fail by several orders of magnitude. Our model contains no parameters adjusted to nucleation simulations, which demonstrates its predictive capability when properly applied and coupled to quantitative thermodynamics and kinetics data. Furthermore, contrary to CNT, the nucleation rate shows a maximum near the spinodal, an effect also observed in diffuse-interface models [29]. We have so far neglected the Tolman correction [77], which accounts for the curvature dependence of surface tension and becomes relevant when the cluster radius is comparable to the interface thickness [78, 79, 80]. While diffuse-interface theories naturally incorporate this correction [21], it must be explicitly added in sharp-interface models. Here, we implement it by multiplying the surface energy contribution in $\Delta\varOmega$ by $(1-2\delta/R)$, retaining only the first order term, where $\delta$ is the conventional Tolman length [77, 81, 82, 21]. Although the curvature correction has only a minor influence on the critical cluster properties, it noticeably affects the nucleation rate due to its exponential dependency in Eq. (2). Nucleation rates with Tolman correction are shown in Fig. 2 for $\delta=0.05\sigma$ and $\delta=0.1\sigma$. To preserve good agreement with MD results, our analysis indicates that the Tolman length must remain small, in line with the value $\delta=0.09\sigma$ reported in Ref. [81] under comparable conditions ($T=0.8\epsilon/k_{\mathrm{B}}$ and a cut-off of $5\sigma$).
Figure 2: Evolution of the steady-state nucleation rate against initial density at $T=0.8\epsilon/k_{\mathrm{B}}$. Our brute-force simulation results are represented by blue up triangles. Results from Diemand et al. [19] are represented by green down triangles. Predictions from our two-variable nucleation theory and CNT are represented by a red solid line and a yellow dotted line, respectively. Predictions from our two-variable nucleation theory incorporated with Tolman correction of $\delta=0.05\sigma$ and $\delta=0.1\sigma$ are represented by the pink and violet thinner solid lines, respectively.
@@
Prompted by the confirmations of its quality, our model can be used to predict the work of formation, see Fig. 3. Tolman correction, that is almost imperceptible on the nucleation barrier, is not shown. At low supersaturation, the three models agree because critical cluster density matches closely the equilibrium liquid density. In the intermediate supersaturation region, both our model and SGA predict much smaller nucleation barriers. It is counter-intuitive at first since the driving force for nucleating a dilute cluster is smaller as compared to a cluster with the equilibrium density, but it is balanced by a much lower surface energy cost. Finally, when approaching the spinodal limit, the two-variable nucleation theory recovers results consistent with diffuse-interface models: the density of critical cluster approaches the initial phase density, its size diverges, and the associated work of formation vanishes. This behavior provides a smooth transition between nucleation and spinodal decomposition, where the energy barrier is zero, as established in diffuse-interface models [23] and recovered in previous nucleation models with two parameters [83, 84, 85, 48, 86, 57, 58, 87, 25, 76, 47]. The surface energy contribution also vanishes at the spinodal limit in our two-variable theory since the cluster density approaches that of the initial phase. CNT, however, fails to capture such effects, highlighting the improved physical accuracy of the two-variable model.
Figure 4: Contour plot of the work of formation overlaid with growth trajectories of critical clusters from simulations at $T=0.8\epsilon/k_{\mathrm{B}}$ and $\rho_{0}=0.03\sigma^{-3}$. The work of formation is computed using our two-variable nucleation theory (without Tolman correction), with the saddle point marked by a red hexagram. The red dash line and blue dotted line indicate the predicted nucleation flux direction and steepest decent direction at the saddle point, respectively. Growth trajectories from four independent simulations are represented by colored crosses, with corresponding mean values shown as dots. Error bars denote standard deviations. $(R,\rho)$ coordinates from theory and simulation are normalized by their corresponding critical values.
@@ Appendix E Incorporation of dynamical density functional theory (DDFT) into nucleation theory
Substituting this expression into Eq. (7), pre-multiplying by $\frac{\partial n}{\partial\boldsymbol{x}}$, and integrating over $r$ from 0 to infinity, we obtain
@@
$\mathbf{D}=D_{\mathrm{diff}}\begin{pmatrix}4\mathrm{\pi}R^{3}\dfrac{(\rho-\rho_{0})^{2}}{\rho_{0}}&4\mathrm{\pi}R^{4}\dfrac{\rho-\rho_{0}}{3\rho_{0}}\\ 4\mathrm{\pi}R^{4}\dfrac{\rho-\rho_{0}}{3\rho_{0}}&\frac{4}{9}\mathrm{\pi}R^{5}\big{(}\frac{1}{5\rho}+\frac{1}{\rho_{0}}\big{)}\end{pmatrix}^{-1}.$ (10)